# Kendall’s $\tau$ Coefficient for Logits Distillation

### v1

@@ 4.3 How RKKD Benefit Optimizing Logits Distillation
In this section, we analyze the optimal solution domain for Kullback-Leibler (KL) divergence and Ranking loss to show how ranking loss affects the optimization process of KL divergence. For KL divergence, the optimal solution domain is achieved when the distribution of the student model aligns perfectly with that of the teacher model, which is equivalent to a linear mapping on the logits.:
$p_{i}^{t}=p_{i}^{s}\quad\forall i\quad\Rightarrow\quad z^{t}_{i}=z^{s}_{i}+c\quad\forall i$ (13)
@@
$\ p_{i}^{t}$ and $\ p_{i}^{s}$ respectively represent the probability outputs of the teacher and the student for the i-th class, and $\ c$ represents any constant number. For the ranking loss, the optimal solution domain is such that for any two indices, the order of the logits output by the teacher and the student is consistent:
$\hskip 5.3ptsgn(z^{t}_{i}-z^{t}_{j})\cdot sgn(z^{s}_{i}-z^{s}_{j})=1\quad\forall i,j$ (14)
@@
which means the optimal solution domain of ranking loss is achieved when there exists a monotonic increasing mapping between the student’s output and the teacher’s output. From the optimal solution domains of the aforementioned two, it is evident that the optimal solution domain for ranking loss is quite lenient and easily attainable, and it encompasses the optimal solution domain for KL divergence. During optimization with the differential form of rank loss, the closer it is to the optimal solution domain of ranking loss, the smaller the gradient provided for optimization becomes. This implies that during model optimization, the ranking loss will prevent the model from falling into the local optimum.
5 Experiment
Datasets and Baselines. In order to validate the efficacy and robustness of our proposed method, we conduct experiments on the CIFAR-100 [16] and ImageNet [24] datasets. CIFAR-100 is a significant dataset for image classification, comprising 100 categories, with 50,000 training images and 10,000 test images, each processed to a size of 32 $\times$ 32. ImageNet is a large-scale dataset utilized for image classification, comprising 1,000 categories, with approximately 1.28 million training images, and 50,000 test images, each processed to a size of 224 $\times$ 224. We apply our RKKD to a variety of logit-based Knowledge Distillation methods, including KD [12], CTKD [20], and MLKD [14]. We also compared various feature-based methods, including FitNet [1], CRD [30], and ReviewKD [4]. In addition, we also extend the proposed RKKD to the transformer student, which has a larger gap with teacher.
Implementation Details. Our experimental setup is primarily based on the settings of the compared baselines [12; 20; 42; 14]. For the CIFAR-100 dataset, we set the batch size to 64, and for the ImageNet dataset, we set the batch size to 512. We employ SGD [28] as the optimizer, with the number of epochs and learning rate settings consistent with the comparative baselines. The hyper-parameters $\alpha$, $\beta$ in Eq. 6 are set to be the same as the compared baselines to maintain fairness, and $\gamma$ are set equal to $\alpha$. We utilize 1 NVIDIA GeForce RTX 4090 to train models on CIFAR-100 and 4 NVIDIA GeForce RTX 4090 for training on ImageNet.
Table 1: CIFAR-100 Heterogeneous Architecture Results. The Top-1 Accuracy (%) is reported as the evaluation metric. The teacher and student have heterogeneous architectures. We incorporate the proposed ranking loss into the existing logit-based methods, with the performance gains indicated in parentheses..
Type Teacher ResNet32$\times$4 ResNet32$\times$4 ResNet32$\times$4 ResNet50 ResNet32$\times$4 WRN-40-2
79.42 79.42 79.42 79.34 79.42 75.61
Student SHN-V2 WRN-16-2 WRN-40-2 MN-V2 SHN-V1 SHN-V1
71.82 73.26 75.61 64.60 70.50 70.50
Feat FitNet [1] 73.54 74.70 77.69 63.16 73.59 73.73
CRD [30] 75.65 75.65 78.15 69.11 75.11 76.05
ReviewKD [4] 77.78 76.11 78.96 69.89 77.45 77.14
Logit KD [12] 74.45 74.9 77.7 67.35 74.07 74.83
KD+RKKD 76.58(+2.13) 75.18(+0.28) 78.50(+0.80) 70.45(+3.10) 75.98(+1.91) 76.13(+1.30)
CTKD [20] 75.37 74.57 77.66 68.67 74.48 75.61
CTKD+RKKD 77.01(+1.64) 75.71(+1.14) 78.61(+0.95) 70.18(+1.51) 76.67(+2.19) 76.80(+1.19)
DKD [42] 77.07 75.7 78.46 70.35 76.35 76.33
DKD+RKKD 77.60(+0.53) 75.99(+0.29) 78.75(+0.29) 70.90(+0.55) 77.36(+1.01) 76.43(+0.10)
MLKD [14] 78.44 76.52 79.26 71.04 77.18 77.44
MLKD+RKKD 78.53(+0.09) 76.83(+0.31) 79.86(+0.60) 71.66(+0.62) 77.63(+0.45) 77.87(+0.43)
Other Settings. In order to prevent the gradient explosions due to too small logit value used in RKKD, we only add normalization to handle the original logit in RKKD. Since using normalization does not affect the ranking, and we do not add normalization to the other loss terms, it does not affect our fairness as a plug-and-play function. More experiments are shown in the Appendix.
5.1 Main Result
CIFAR-100 Results. In our study, we conducted a comparative analysis of Knowledge Distillation (KD) outcomes across various teacher-student [9; 26; 39; 41; 13; 25] configurations as shown in Tab. 1 and Tab. 2. While Tab. 1 presents cases where the teacher and student models share the heterogeneous architecture, Tab. 2 illustrates instances of homogenous structures. Furthermore, as a plug-and-play method, we applied our RKKD in multiple logit-based distillation techniques. The incorporation of RKKD resulted in an average improvement of 1.83% in vanilla KD. In the context of existing state-of-the-art (SOTA) logits-based methods, both CTKD and MLKD exhibited significant gains. Notably, a subset of students surpassed their respective teachers after the distillation.
Table 2: CIFAR-100 Homogenous Architecture Results. The Top-1 Accuracy (%) is reported as the evaluation metric. The teacher and student have homogeneous architectures. We incorporate the proposed ranking loss into the existing logit-based methods, with the performance gains indicated in parentheses.
Type Teacher ResNet32$\times$4 VGG13 WRN-40-2 ResNet110
@@ 5.1 Main Result
Feat FitNet [1] 73.50 71.02 72.24 68.99
CRD [30] 75.51 73.94 74.14 71.46
ReviewKD [4] 75.63 74.84 75.09 71.34
Logit KD [12] 73.33 72.98 73.54 70.67
KD+RKKD 74.74(+1.41) 74.14(+1.16) 74.49(+0.95) 71.09(+0.42)
CTKD [20] 73.39 73.52 73.93 70.99
CTKD+RKKD 75.59(+2.2) 74.76(+1.24) 74.86(+0.93) 71.08(+0.09)
DKD [42] 76.32 74.68 74.81 71.06
DKD+RKKD 76.61(+0.29) 75.1(+0.42) 74.94 (+0.13) 71.84(+0.78)
MLKD [14] 77.08 75.18 75.35 71.89
MLKD+RKKD 77.25(+0.17) 75.35(+0.17) 76.08(+0.73) 72.35(+0.46)
@@
ImageNet Results. The results on ImageNet of KD in terms of top-1 and top-5 accuracy are compared in Tab.3. Our RKKD can achieve consistent improvement on the large-scale dataset as well.
@@ 5.2 Extensions
KD for Transformer. Further exploration was conducted on the implications of rank loss when employing a transformer as both the teacher and the backbone. The experimental results indicate that the incorporation of RKKD yields significant improvements over the conventional knowledge distillation, as shown in Tab. 4. This denotes that RKKD is also applicable to distillation challenges predicated on transformer architectures. The implementation details of the Transformer experiments are attached in the Appendix.
Table 4: KD for Transformer. The Top-1 Accuracy (%) on the validation set of CIFAR-100. The Teacher model is ResNet56.
Student DeiT-Tiny [31] T2T-ViT-7 [37] PiT-Tiny [11] PVT-Tiny [33]
@@
KD+RKKD 73.25(+2.14) 72.49(+0.77) 74.33(+0.30) 73.42(+0.96)
KD for Object Detection. To further verify the generality of our proposed method in downstream tasks, we also conduct experiments under the setting of object detection. The results show that our proposed ranking loss can improve the performance of knowledge distillation method in object detection and achieve better performance than the same period of the feature-based object detection method, which is shown in Tab. 5.
Table 5: KD for Object Detection. All experiments are conducted on COCO2017 with the teacher as ResNet50 and the student as MobileNet-V2.
Method AP AP50 AP70
KD 30.13 50.28 31.35
FitNet[[1]] 30.20 49.80 31.69
FGFI[[32]] 31.16 50.68 32.92
KD+RKKD 31.99(+21.86) 53.80(+3.52) 33.37(+2.02)
Different Forms of Ranking Loss. In our investigation, we examined the performance of the three forms of distillation presented in Sec. 4.2, as depicted in Tab. 6. Form1 exhibited the most superior performance, followed by Form2, with Form3 trailing yet still enhancing the efficacy of the original knowledge distillation. This suggests that within the ranking loss, it is imperative to optimize considering the magnitude of the teacher’s logits differences as a coefficient for the loss, rather than merely optimizing as a sign function.
Loss Form KD KD+Form1 KD+Form2 KD+Form3
Similar Structure 72.01 73.47(+1.46) 73.42(+1.41) 73.38(+1.37)
Different Structure 72.74 73.50(+0.76) 73.36(+0.62) 73.05(+0.31)
Ablation Study. We conducted extensive ablation studies to investigate the effectiveness of RKKD under various settings of $k$ and coefficients. Tab. 8 presents the distillation outcomes across different $k$ configurations. It is observed that a larger $k$ value yields superior performance, suggesting that the differential form of ranking loss more closely approximates the Kendall $\tau$ coefficient, thereby imparting stronger ranking knowledge. Tab. 7 delineates the performance of RKKD under varying coefficients. This setting effectively aligns the classification outcomes of the teacher and student models, thus significantly enhancing the distillation effect.
@@ 5.3 Analysis
Figure 3: Loss Landscape. The Left landscape shows the suboptimal solutions in the distillation task. After adding our ranking loss, the suboptimal solution is significantly reduced, as shown on the Right.
Visualization of Loss Landscape. To further investigate the role of ranking loss in the distillation process, we visualized the loss landscapes [19] of student models with and without the application of ranking loss during Knowledge Distillation (KD), as depicted in Fig. 4. It is evident that the student models distilled with ranking loss exhibit a markedly flatter loss landscape and fewer local optima compared to those without it. We hypothesize that during the optimization process, ranking loss can filter out certain local optima that, despite presenting a better overall loss performance, yield poorer classification outcomes. Consequently, ranking loss can effectively enhance the generalization performance of student models.
Figure 4: RKKD using Top-K and Min-K channels. Top-1 Accuracy (%) of Top-k/Min-k Ranking Loss. 0 and 100 on the x-axis represent KD and KD+Rank, respectively.
Top-k & Min-k Ranking. To further validate the beneficial knowledge present in smaller channels, we conducted comparative experiments using the top 10%, top 30%, top 50%, and min 10%, min 30%, min 50% channels. The experiments were performed across four combinations of homogeneous and heterogeneous teacher-student pairs, and the results are presented in Fig. 4. We observed that using the min-k channels achieves results similar to those obtained with top-k channels (as shown in Fig. 4(a),(b), and (c)). Additionally, in some cases, min-k channels provide even more beneficial information to aid student learning (as demonstrated in Fig. 4(d)), which indicates that the smaller channels also contain rich knowledge.

### v2

@@ 4.2.1 From the Perspective of Gradient: Ranking Loss Cares about Low-Probability Channels
In this section, we analyze the optimal solution domain for Kullback-Leibler (KL) divergence and ranking loss to show how ranking loss affects the optimization process of KL divergence. For KL divergence, the optimal solution domain is achieved when the distribution of the student model aligns perfectly with that of the teacher model, which is equivalent to a linear mapping on the logits:
$q_{i}^{t}=q_{i}^{s}\quad\forall i\quad\Rightarrow\quad z^{t}_{i}=z^{s}_{i}+c\quad\forall i$ (11)
@@
$\ q_{i}^{t}$ and $\ q_{i}^{s}$ respectively represent the probability outputs of the teacher and the student for the i-th class, and $\ c$ represents any constant number. For the ranking loss, the optimal solution domain is such that for any two indices, the order of the logits output by the teacher and the student is consistent:
$\hskip 5.3pt\mathrm{sgn}(z^{t}_{i}-z^{t}_{j})\cdot\mathrm{sgn}(z^{s}_{i}-z^{s}_{j})=1\quad\forall i,j$ (12)
@@
It implies that the optimal solution domain for ranking loss is quite lenient and easily attainable, encompassing the optimal solution domain for KL divergence. Therefore, in the later stage of training, the ranking loss will not hinder the optimization of KL divergence, which means our ranking loss can perfectly be used as an auxiliary loss.
Table 2: CIFAR-100 Homogenous Architecture Results. The Top-1 Accuracy (%) is reported as the evaluation metric. The teacher and student have homogeneous architectures. We incorporate the proposed ranking loss into the existing logit-based methods, with the performance gains indicated in parentheses. The best and second best results are emphasized in bold and underlined.
KD Teacher ResNet32$\times$4 VGG13 WRN-40-2 ResNet110
@@
FitNet (Adriana et al., 2015) 73.50 71.02 72.24 68.99
CRD (Tian et al., 2019) 75.51 73.94 74.14 71.46
ReviewKD (Chen et al., 2021) 75.63 74.84 75.09 71.34
DIST (Huang et al., 2022) 76.31 73.80 74.73 71.40
LSKD (Sun et al., 2024) 78.28 75.22 75.56 72.27
KD (Hinton et al., 2015) 73.33 72.98 73.54 70.67
KD+Ours 74.74(+1.41) 74.14(+1.16) 74.49(+0.95) 71.09(+0.42)
CTKD (Li et al., 2023) 73.39 73.52 73.93 70.99
CTKD+Ours 75.59(+2.2) 74.76(+1.24) 74.86(+0.93) 71.08(+0.09)
DKD (Zhao et al., 2022) 76.32 74.68 74.81 71.06
DKD+Ours 76.61(+0.29) 75.1(+0.42) 74.94 (+0.13) 71.84(+0.78)
MLKD (Jin et al., 2023) 77.08 75.18 75.35 71.89
MLKD+Ours 77.25(+0.17) 75.35(+0.17) 76.08(+0.73) 72.35(+0.46)
4.2.3 From the Perspective of Classification: Ranking Loss Helps Student Classify Correctly
As illustrated in Fig. 1, using KL loss may lead to a smaller loss yet result in incorrect classification. The tendency to fall into local optima will hinder the optimization process in distillation. Although the use of cross-entropy loss can mitigate this issue, it does not incorporate information from the teacher model. Consequently, in experiments, cross-entropy loss is often assigned a very small weight, which also leads to a small gradient. For instance, KD (Hinton et al., 2015)$\mathcal{L}_{\text{CE}}:\mathcal{L}_{\text{KL}}=0.1:0.9$; DKD (Zhao et al., 2022) $\mathcal{L}_{\text{CE}}:\mathcal{L}_{\text{TCKL}}:\mathcal{L}_{\text{NCKL}}=1:1:8$; MLKD (Jin et al., 2023) $\mathcal{L}_{\text{CE}}:\mathcal{L}_{\text{KL}}=0.1:9$; LSKD (Sun et al., 2024) $\mathcal{L}_{\text{CE}}:\mathcal{L}_{\text{KL}}=0.1:9$. As a result, cases with lower KL loss but incorrect classification still frequently occur. In contrast, by aligning the rank between student and teacher, ranking loss helps KL divergence avoid such suboptimal situations. Also, the ranking loss can incorporate information from the teacher model. In this way, ranking loss helps students classify correctly, avoiding the suboptimal case in logit distillation. Furthermore, a model with better generalization should have a more reasonable rank of logits. For instance, recognizing that a tiger is more similar to a cat than to a fish. By learning such inter-class relationships, the student can improve classification performance and enhance its representational capacity.
5 Experiments
Datasets. In order to validate the efficacy and robustness of our proposed method, we conduct extensive experiments on three datasets. 1) CIFAR-100 (Krizhevsky et al., 2009) is a significant dataset for image classification, comprising 100 categories, with 50,000 training images and 10,000 test images. 2) ImageNet (Russakovsky et al., 2015) is a large-scale dataset utilized for image classification, comprising 1,000 categories, with approximately 1.28 million training images and 50,000 test images. 3) MS-COCO (Lin et al., 2014) is a mainstream dataset for object detection comprising 80 categories, with 118,000 training images and 5,000 test images.
Baselines. As a plug-and-play loss, we apply the proposed ranking loss to various logit-based methods, including KD (Hinton et al., 2015), CTKD (Li et al., 2023), DKD (Zhao et al., 2022), and MLKD (Jin et al., 2023), to verify whether it can bring performance gains to knowledge distillation methods. We also compare it with various feature-based methods, including FitNet (Adriana et al., 2015), CRD (Tian et al., 2019), and ReviewKD (Chen et al., 2021). Additionally, we compare it with other auxiliary losses and modified KL divergence methods, including DIST (Huang et al., 2022) and LSKD (Sun et al., 2024).
Implementation Details. To ensure the robustness of the proposed plug-and-play loss without introducing excessive configurations, we maintain the same experimental settings as the baselines used (KD+Ours and KD share the same experimental setups for example). We set the batch size to 64 for CIFAR-100, 512 for ImageNet and 8 for COCO. We employ SGD (Sutskever et al., 2013) as the optimizer, with the number of epochs and learning rate settings consistent with the comparative baselines. The hyper-parameters $\alpha$, $\beta$ in Eq. 6 are set to be the same as the compared baselines to maintain fairness, and $\gamma$ are set equal to $\alpha$. We utilize 1 NVIDIA GeForce RTX 4090 to train models on CIFAR-100 and 4 NVIDIA GeForce RTX 4090 for training on ImageNet. The algorithm of our method can be found in Appendix A.4.
Other Settings. In order to prevent the occasional gradient explosions at the start of optimization due to the too variant logits caused by random initialization of student, we add and only add normalization in ranking loss, which does not affect our fairness as a plug-and-play function. We conduct experiments about ablation of normalization in Appendix A.2, and the algorithm can be found in Appendix A.4.
5.1 Main Result
CIFAR-100 Results. In our study, we conduct a comparative analysis of Knowledge Distillation (KD) outcomes across various teacher-student (He et al., 2016; Simonyan & Zisserman, 2014; Zagoruyko & Komodakis, 2016; Zhang et al., 2018; Howard et al., 2017; Sandler et al., 2018) configurations. While Tab. 1 presents cases where the teacher and student models share the heterogeneous architecture, Tab. 2 illustrates instances of homogenous structures. Furthermore, as a plug-and-play method, we applied ranking loss in multiple logit-based distillation techniques. The incorporation of ranking loss resulted in an average improvement of 1.83% in vanilla KD. In the context of existing state-of-the-art (SOTA) logits-based methods, including DKD, CTKD, and MLKD, significant gains are made.
@@ 5.1 Main Result
ImageNet Results. The results on ImageNet of KD in terms of top-1 and top-5 accuracy are compared in Tab. 3. Our proposed method can achieve consistent improvement on the large-scale dataset as well. Additional ImageNet experiments are provided in Appendix A.9
@@ 5.2 Extensions
Knowledge Distillation for Transformer. To validate the effectiveness of our plug-and-play ranking loss on ViT and to assess its performance when facing larger teacher-student structural differences, we conduct experiments using ViT students. The experimental results indicate that the incorporation of ranking loss yields significant improvements over the conventional knowledge distillation, as shown in Tab. 4. This denotes that our proposed method is also applicable to distillation challenges predicated on transformer architectures. The implementation details of the Transformer experiments are attached in the Appendix.
Table 4: Knowledge Distillation for Transformer. The Top-1 Accuracy (%) on the validation set of CIFAR-100. The Teacher model is ResNet56.
Student DeiT-Tiny T2T-ViT-7 PiT-Tiny PVT-Tiny
@@
KD+Ours 73.25(+2.14) 72.49(+0.77) 74.33(+0.30) 73.42(+0.96)
Table 5: Knowledge Distillation for Object Detection. All experiments are conducted on COCO2017 with the teacher as ResNet50 and the student as MobileNet-V2.
Method AP AP50 AP75
KD (Hinton et al., 2015) 30.13 50.28 31.35
FitNet (Adriana et al., 2015) 30.20 49.80 31.69
FGFI (Wang et al., 2019) 31.16 50.68 32.92
CTKD (Li et al., 2023) 31.39 52.34 31.35
KD+Ours 31.99(+1.86) 53.80(+3.52) 33.37(+2.02)
Knowledge Distillation for Object Detection. To further verify the generality of our proposed method in downstream tasks, we also conduct experiments under the setting of object detection. The results show that our proposed ranking loss can improve the performance of knowledge distillation method in object detection and achieve better performance than the same period of the feature-based object detection method, which is shown in Tab. 5.
Ablation of Weight and Steepness Parameter. We conduct extensive ablation studies to investigate the effectiveness of ranking loss under various settings of $k$ and coefficients. Tab. 7 presents the distillation outcomes across different $k$ configurations. It is observed that a larger $k$ value yields superior performance, suggesting that the differential form of ranking loss more closely approximates the Kendall $\tau$ coefficient, thereby imparting stronger ranking knowledge. Tab. 6 delineates the performance of ranking loss under varying coefficients. This setting effectively aligns the classification outcomes of the teacher and student models, thus significantly enhancing the distillation effect.
@@
Ablation of Temperature. There has been extensive and detailed research on the temperature of the KL loss (e.g., CTKD(Li et al., 2023), MLKD(Jin et al., 2023), LSKD(Sun et al., 2024)), achieving significant results. Unlike these studies, our approach aims to move away from focusing on the KL divergence and instead explore plug-and-play auxiliary losses to guide the KL divergence. Therefore, in our experiments, the temperature and other parameters of the KL divergence are kept the same as those in the respective baselines (with dynamic/multi-level temperatures used in CTKD+Ours and MLKD+Ours). The proposed ranking loss does not have a temperature parameter because the temperature does not affect the order of logits. Instead, the control over the distribution can be achieved by manipulating the steepness parameter $k$ of the sign function, and ablation experiments for $k$ are shown in Tab. 7. Meanwhile, although the temperature helps control the logit distribution, additional measures are needed to prevent the KL divergence from falling into suboptimal. We conducted an ablation study on the temperature, as shown in Tab. 8. The results demonstrate that ranking loss as a plug-and-play loss can bring further improvements under multiple temperature settings.
Figure 4: Ranking Loss using Top-K and Min-K channels. Top-1 Accuracy (%) of Top-k/Min-k Ranking Loss. 0 and 100 on the x-axis represent KD and KD+Rank, respectively.
Comparison between Ranking loss and KL Divergence. We separately evaluate distillation with KL divergence alone and with the proposed ranking loss alone; the results are presented in Tab. 9. We observe that using the ranking loss alone achieves performance comparable to using KL divergence alone, and that combining the two objectives yields further gains. These findings substantiate the efficacy of the proposed ranking loss.
Temperature T = 4 T = 5 T = 6 T = 10
KD 73.33 73.39 73.43 73.55
KD+Ours 73.56(+0.23) 73.49(+0.10) 74.36(+0.93) 74.04(+0.49)
Table 9: Experiments on Combining DIST.
Teacher ResNet32$\times$4 ResNet32$\times$4 WRN-40-2 VGG13 WRN-40-2
Student SHN-V1 WRN-40-2 SHN-V1 VGG8 WRN-40-1
KD (Only KL) 74.07 77.70 74.83 72.98 73.54
Only Ranking Loss 75.60 78.04 75.81 73.72 74.00
KL+Ranking Loss 75.98 78.50 76.13 74.14 74.49
More Experiments. Additional experiments and discussions, including Derivation of the KL Loss and Ranking Loss, Ablation of Normalization, and Different Forms of Ranking Loss, are provided in the Appendix. Please refer to the Sec. A for further details.
@@ 5.3 Analysis
Accuracy & Loss Curves with Ranking Loss. The accuracy and loss curves of KD and KD+Ours, as shown in Fig. 5, demonstrate how ranking loss aids in optimization. The middle figure shows that the precise alignment of KL divergence also makes channel ranking more ordered, but adding ranking loss achieves a more consistent ranking more quickly. The right figure shows that in the early stages, ranking loss accelerates the reduction of KL loss and reduces its oscillation in suboptimal regions. In the later stages, ranking loss does not interfere with KL divergence and ultimately reaches a better position. The left figure shows that with ranking loss, student achieves leading accuracy at all stages. This indicates that the addition of the ranking loss helps the model converge and achieve better generalization and performance.
Figure 5: Accuracy & Loss Curves with Ranking Loss. Left: Top-1 Test Accuracy (%) Curve. Middle: Loss Curve of Ranking Loss. Right: Loss Curve of KL Divergence.
Figure 6: Loss Landscape. Left: The left landscape shows the suboptimal solutions in the distillation task. Right: After adding our ranking loss, the suboptimal solution is significantly reduced, as shown on the right.
Visualization of Loss Landscape. To further investigate the role of ranking loss in the distillation process, we visualized the loss landscapes (Li et al., 2018) of student models with and without the application of ranking loss during Knowledge Distillation (KD), as depicted in Fig. 4. It is evident that the student models distilled with ranking loss exhibit a markedly flatter loss landscape and fewer local optima compared to those without it. We hypothesize that during the optimization process, ranking loss can filter out certain local optima that, despite presenting a better overall loss performance, yield poorer classification outcomes. Consequently, ranking loss can effectively enhance the generalization performance of student models.
Top-k & Min-k Ranking. To further validate the beneficial knowledge present in low-probability channels, we conducted comparative experiments using the top 10%, top 30%, top 50%, and min 10%, min 30%, min 50% channels. The experiments were performed across four combinations of homogeneous and heterogeneous teacher-student pairs, and the results are presented in Fig. 4. We observed that using the min-k channels achieves results similar to those obtained with top-k channels (as shown in Fig. 4(a),(b), and (c)). Additionally, in some cases, min-k channels provide even more beneficial information to aid student learning (as demonstrated in Fig. 4(d)), which indicates that the low-probability channels also contain rich knowledge.
Figure 7: Ranking Alignment Comparison. Left: KD. Right: KD+Ours.
Further visualization of Ranking. We show in Fig. 3 that ranking loss balances the focus on high-probability and low-probability channels. To further illustrate the role of ranking loss, we present a visualization of the ranking results generated by KD and KD+Ours in Fig. 7. It can be seen that after adding ranking loss, the channels are more ordered overall, which is more obvious in the low-probability channel part. This is because KL divergence does not pay attention to sorting and neglects low-probability channels. Although the large channels in KD receive more attention, the alignment value alone cannot achieve order consistency, thus losing some knowledge. The results indicate that our approach consistently enhances channel alignment, underscoring the robustness and general applicability of our method, particularly in its focused attention on low-probability channels.
Further Ablation Study of the Hyperparameters. To further substantiate the generalization capability and robustness of our method, we conducted comprehensive ablation studies using different combinations of coefficients. As illustrated in Fig. 8, our method consistently maintains high performance across various settings, underscoring the strong generalization ability and universal applicability of our method.
Figure 8: Sensitivity Analysis. Left: Sensitivity of $\alpha-\gamma$. Right: Sensitivity of $\beta-\gamma$.